The content of a sphere with a volume of 324 cm^3 is poured into a cylindrical vessel of the same radius as that of the sphere. If the cylindrical vessel is only 3/4 filled, find the height of the cylindrical vessel. Use pi = 3.14

Answer by mananth(16946)

The content of a sphere with a volume of 324 cm^3 is poured into a cylindrical vessel of the same radius as that of the sphere. If the cylindrical vessel is only 3/4 filled, find the height of the cylindrical vessel. Use pi = 3.14

Volume of sphere = 4/3 pi r^3= 324
r^3= 3*324/(4 pi)
r^3 = 243/pi
r^3=77.39
r=4.26
Volume of fill = pi r^2 h
324 = pi *r^2 *h
324/(pi* r^2) =h
324/(3.14*(4.26)^2)=h
h=5.68 cm